How much pure water must be mixed with 4 pints of 50% acetic acid to produce a mixture that is 33% acetic acid?


Answer by mananth(16949)

------------percent----------------quantity
solution------50---------------- 4
water---------0 ---------------- x
Total--------33 ---------------- 4+x
...
50*4+0*x=33(4+x)

200=132 +33x

-33x=132-200

-33x=-68
/-33
x=2.06 liters of water